Math & Logic · Ages 12–15
Riddle #152
If you flip a fair coin 3 times, what is the probability of getting at least one heads?
You solved this one before
Answer: 7/8
1 − (probability of all tails) = 1 − (1/2)³ = 1 − 1/8 = 7/8.
Math & Logic · Ages 12–15
If you flip a fair coin 3 times, what is the probability of getting at least one heads?
You solved this one before
Answer: 7/8
1 − (probability of all tails) = 1 − (1/2)³ = 1 − 1/8 = 7/8.